If 2x+7y=17 and 2x+5y=11, what is the value of 2y?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

If 2x+7y=17
and 2x+5y=11,
what is the value of 2y?
------------------------------
Subtract the 2nd equation from the 1st to get:
2y = 6
============
Cheers,
Stan H.

Answer by richwmiller(17219)


We can solve for a numerical value or a value in terms of x.
2x+7y=17 and 2x+5y=11
But notice if we subtract the second equation from the first we get
2x+7y=17
2x+5y=11
2y=6